The Treasury Department auctioned $21 billion in three-month bills in denominations of ten thousand dollars at a
discount rate of 4.965%, according to the March 13, 2007, issue of the Chicago Sun-Times. What would be the effective rate of interest? Round your answer to the nearest hundredth percent.

Answer by stanbon(75887)

The Treasury Department auctioned $21 billion in three-month bills in denominations of ten thousand dollars at a
discount rate of 4.965%, according to the March 13, 2007, issue of the Chicago Sun-Times. What would be the effective rate of interest? Round your answer to the nearest hundredth percent.
------
amt. borrowed:: 21 billion
amt. actually received 0.95035*21 billion = 19.95735 billion
interest paid in 3 months = 1.04265 billion
----
I = P*r*t
1.04265 = 19.95735*r*(3/12)
----
rate = 0.209 = 20.9%
--------------
Cheers,
Stan H.
-----------